OpenStudy (zaxoanl):

how to evaluate the integrals x/(3+sqrt(x)) improper integral from -x to 0 e^xdx

OpenStudy (anonymous):

\[\int\limits_{}^{} \frac{ x }{ 3+\sqrt x }dx\]

OpenStudy (anonymous):

is this your question?

OpenStudy (zaxoanl):

yes

OpenStudy (zaxoanl):

i let U=\[\sqrt{x}\] U^2 = x 2Udu=dx

OpenStudy (anonymous):

Make u=x^2

OpenStudy (anonymous):

du = 2 x dx

OpenStudy (zaxoanl):

then im stuck at \[\int\limits_{?}^{?} 2u^3/(3+u)du\]

OpenStudy (zaxoanl):

oh ok let me try that

OpenStudy (anonymous):

I am changing my mind. Wait a minute

OpenStudy (zaxoanl):

sorry but im stuck at \[\sqrt{x}\] when substituting in the u value in

OpenStudy (anonymous):

Actually, your first substitution works better, I think

OpenStudy (anonymous):

Notice that \[ \frac{2 u^3}{u+3}=2 u^2-6 u-\frac{54}{u+3}+18 \]

OpenStudy (anonymous):

By doing long division. The rest is easy

OpenStudy (zaxoanl):

thank you i get it now

OpenStudy (anonymous):

Your final answer should look like \[ \frac{2 x^{3/2}}{3}-3 x+18 \sqrt{x}-54 \log \left(\sqrt{x}+3\right) \]

OpenStudy (anonymous):

YW

OpenStudy (zaxoanl):

can you help me on the improper integral

OpenStudy (zaxoanl):

\[\int\limits_{-\infty}^{0} e^xdx\]

OpenStudy (hossamhoussien):

Use this calculator to understand the problem you stuck at NOT for solving assignments or something like that and remember that it's all about understanding Good luck ;) https://www.symbolab.com/solver/integral-calculator

OpenStudy (priyar):

do u know what is e^x integration?

OpenStudy (hossamhoussien):

if it was indefinite integral it's simply \[e^x+c\] if it was definite integral it's simply\[\int\limits_{Lower } ^ {Upper} e^x dx =e^{Upper} - e^{Lower} \]

OpenStudy (mathmale):

@zaxoanl: \[\int\limits\limits_{-\infty}^{0} e^xdx\] ... 1) is an improper integral due to the negative infinity lower limit. 2) is a definite integral. See @HossamHoussien 's "definite integral" equation, above. Applying this definite integral formula to your \[\int\limits\limits\limits_{-\infty}^{0} e^xdx, \] \[\int\limits\limits\limits_{-\infty}^{0} e^xdx\]

OpenStudy (mathmale):

\[\int\limits\limits\limits\limits_{-\infty}^{0} e^xdx=e^0-limit.as.x \rightarrow negative.infinity.of.e^x\]

OpenStudy (mathmale):

Can you finish evaluating this definite integral?

OpenStudy (anonymous):

\[e^0-e ^{-\infty }=1-\frac{ 1 }{ e^\infty }=1-0=1\]